# How informative are summaries of the cosmic 21-cm signal?

### v1

@@ 2.2 Finite differencing
In order to numerically evaluate Eq. 4, we group the terms into ?fixed? ($\bm{\mu}$, $\bm{\Sigma}$, $\bm{\Sigma}^{-1}$) and ?differentiated? ($\nabla_{\bm{\theta}}\,\bm{\mu}$, $\nabla_{\bm{\theta}}\,\bm{\Sigma}$). We calculate ?fixed? terms from realizations of the summary around the same parameter value $\bm{\theta}^{*}$. To compute the mean, we can use the maximum likelihood estimator (MLE):
@@
where the index $i$ labels different realizations of initial conditions and other sources of stochasticity (e.g. telescope noise). It is known that such an estimator of the mean is unbiased. However for the covariance matrix the situation is more complex, as the MLE
@@
is biased. One can show that the estimators of the covariance and its inverse can be unbiased in the following way (e.g. Hartlap et al. 2007):
@@
A similar approach is used to estimate the ?differentiated parameters?. In the case in which the simulator is not differentiable, we can use finite differencing to estimate gradients of the mean and covariance:
@@
Here, $\delta\bm{\theta}_{m}$ represents a small change in the parameter over which the derivative is computed and $\delta\bm{d}(\bm{\theta}^{*})_{i}/\delta\bm{\theta}_{m}$ corresponds to the numerical derivative of a given summary realization (i.e. fixing all sources of stochasticity).
@@ 2.3 The distribution of the Fisher information over parameter space
Below we will show examples of common 21-cm summaries evaluated at a ?fiducial? set of parameters, $\bm{\theta}_{\text{fid}}$. Unfortunately, the cosmic 21-cm signal does not yet have a well-defined fiducial model and our choice of $\bm{\theta}_{\text{fid}}$ is only a ?best guess?. Therefore in order to properly assess the quality of a particular summary, we sample many points $\bm{\theta}^{*}$ from the prior $P(\bm{\theta})$, calculating the Fisher information $\det\bm{F}(\bm{\theta}^{*})$ at each sample. We use this prior-weighted distribution of Fisher information as the main metric for assessing the quality of a given summary.
3 The cosmic 21-cm signal and its summary statistics
Figure 1: Schematic illustrating our pipeline for producing data summaries for a given astrophysical parameter combination (adapted from Prelogović & Mesinger 2023). Starting from a realization of the cosmological signal simulated with 21cmFAST, we remove the mean at each redshift and add a telescope noise realization corresponding to a 1000h observation with SKA1-Low. The resulting lightcone is compressed into different summaries. Explicitly-defined summaries are marked in blue, while neural network summaries are marked in red. This realization is computed at the fiducial parameter set, which used only for IMMN training and visualizations of the summaries. See text for more details.
For a given sample of astrophysical parameters, $\bm{\theta}^{*}$, we compute a realization of mock data using the following steps (illustrated in Fig. 1):
@@ 3 The cosmic 21-cm signal and its summary statistics
    Sampling cosmological initial conditions, we simulate a realization of the 21-cm lightcone using the 21cmFAST code.
@@
    We subtract the mean from each frequency slice, mimicking an interferometric observation.
@@
    We simulate a $1000\mathrm{h}$ tracked scan with the SKA1-low telescope, including the corresponding $uv$ coverage and realization of the instrument noise.
@@
    We compress the resulting lightcone into a given summary observation.
We summarize these steps in more detail below.
@@
The cosmic 21-cm signal is simulated using the semi-numerical code, 21cmFAST v3555https://github.com/21cmfast/21cmFAST (Mesinger & Furlanetto, 2007; Mesinger et al., 2011; Murray et al., 2020). In brief, we use the Park et al. (2019) galaxy model, varying five parameters that characterize the unknown UV and X-ray properties of high-$z$ galaxies through scaling relations:
@@
  • $\bm{f_{*,10}}$ the fraction of galactic gas in stars, normalized at the halo mass of $10^{10}M_{\odot}$. The stellar to halo mass relation of the faint galaxies that drive reionization is well described by a power law: $M_{*}/M_{h}=f_{*,10}(M_{h}/10^{10}M_{\odot})^{0.5}(\Omega_{b}/\Omega_{m})$ (see Park et al. 2019 and references therein).
@@
  • $\bm{f_{\text{esc},10}}$ the ionizing UV escape fraction, normalized at the halo mass $10^{10}M_{\odot}$. Again, the (mean) escape fraction is assumed to be a power law with halo mass, and here we fix its index to $\alpha_{\text{esc}}=-0.5$ (e.g. Qin et al. in prep).
@@
  • $\bm{M_{\text{turn}}}$ the characteristic host halo mass below which galaxies are inefficient at forming stars, due to slow gas accretion, supernovae feedback, and/or photoheating.
@@
  • $\bm{E_{0}}$ the characteristic X-ray energy below which photons are absorbed by the interstellar medium (ISM) of the host galaxy.
@@
  • $\bm{L}_{\bm{X}<2\mathrm{keV}}/\mathrm{SFR}$ the X-ray soft band (in the energy range $E_{0}$ – 2 keV) luminosity per star formation rate (SFR), in units of $\mathrm{erg\,s^{-1}\,keV^{-1}\,M_{\odot}^{-1}\,yr}$. The SFR is taken to be: $\dot{M}_{*}=M_{*}/(0.5H(z)^{-1})$, where the Hubble time, $H(z)^{-1}$, also scales with the dynamical time at the virial radius.
@@
Our fiducial, $\bm{\theta}_{\text{fid}}$ parameter vector corresponds to the following: $\log_{10}f_{*,10}=-1.3$, $\log_{10}f_{\text{esc}}=-1$, $\log_{10}M_{\text{turn}}[M_{\odot}]=8.7$, $\log_{10}L_{X<2\mathrm{keV}}/\text{SFR}[\mathrm{erg\,s^{-1}\,keV^{-1}\,M_{\odot}^{-1}\,yr}]=40$, $E_{0}[\mathrm{keV}]=0.5$. This particular choice is only used when training the IMMN, and for the visualizations of the summaries below. For more details on the model and its motivations, see Park et al. 2019 and references therein.
@@
To simulate the observation we follow the procedure described in Prelogović & Mesinger (2023) (c.f. Fig. 1). In brief, we calculate the 21-cm signal from redshifts 30 to 5, on a 300 Mpc coeval cube, with 1.5 Mpc resolution. The snapshots of the evolved coeval cube are then stacked to produce the final lightcone of the signal. After the subtraction of the mean for every sky-plane slice, we add thermal noise corresponding to a a $6\,\mathrm{h/day}$, $10\,\mathrm{s}$ integration time tracked scan with SKA1-Low for a $1000\,\mathrm{h}$ observation in total. Before computing each summary, we additionally smooth the lightone with a box-car filter obtaining a final resolution of 6 Mpc. This final step does not impact our results as it is comparable to the SKA1-Low beam, but is needed in order to minimize GPU memory usage for certain summaries (see below for more details). We also note that in this work we make the optimistic assumption of perfect foreground removal, thus exploring the maximum future potential of our set of summary statistics.
@@
The resulting 3D lightcone is compressed into a given summary statistic. We explore the following summaries in this work:
@@
    - spherically-averaged 1D power spectrum,
@@
    - cylindrically-averaged 2D power spectrum,
@@
    - 2D wavelet scattering transform,
@@
    - recurrent neural network,
@@
    - information maximizing neural network.
@@
    - concatenation of the 2DPS and IMNN summaries.
The first four are explicitly defined while the following two are ?learned? by neural networks. The last, 2DPS + IMNN, is a combination of the two individual summaries. We describe them in detail below.
3.1 Spherically-averaged (1D) Power Spectrum
Figure 2: 1D power spectrum for the fiducial lightcone shown in Figure 1. Different colors correspond to different wavemodes while solid (dashed) correspond to with (without) noise. Comparing the solid and dashed lines, we see that the signal is noise dominated at higher $k$ modes and redshifts (see also Fig. 1). All power spectra are computed from comoving cubes extracted from the lightcone, centered on the redshifts indicated by the vertical dashed lines.
As discussed in the introduction, the spherically-averaged (1D) power spectrum, 1DPS, is the most common choice of a summary statistic. Despite the cosmic 21-cm signal being non-Gaussian, the 1DPS is a natural choice when seeking to maximize the S/N of an interferometric first detection.
The 1DPS of the 21-cm signal $\delta T_{b}(\bm{x},z)$, where $\bm{x}$ is the sky-plane coordinate and $z$ the redshift, is defined as:
@@ 3.1 Spherically-averaged (1D) Power Spectrum
[…]
@@ 3.2 Cylindrical (2D) Power Spectrum
Figure 3: 2D power spectrum of the fiducial model shown in Figure 1. Different columns show the 2DPS for the pure cosmic signal and including thermal noise, as labeled on the top. Each pair corresponds to a different central redshift, coinciding with those used in the 1DPS summary (see Fig. 2). Furthermore, the bins between 1DPS and 2DPS coincide as close as possible. As in the 1DPS case, one can notice that high-$k$ modes and high redshifts are noise dominated.
The cylindrically-averaged 2D PS distinguishes between sky-plane ($k_{\perp}$) and line-of-sight ($k_{\|}$) modes. Redshift space distortions (e.g. Bharadwaj & Ali 2005; Barkana & Loeb 2005) and lightcone evolution (e.g. Greig & Mesinger 2018; Mondal et al. 2018) result in an anisotropic cosmic signal. Therefore, the 2DPS should encode more physical information compared to the 1DPS. Moreover, the instrument/foregrounds are better characterized in the 2D PS (e.g. see the review in Liu et al. 2014b), making it a natural summary observable for preliminary, low S/N measurements.
The 2DPS is defined as:
@@
where the Fourier transform is performed on the same lightcone chunks as in the 1D case, with the difference being that the expectation value is calculated over $(k_{\perp},k_{\|})$ bins. Here we pick $3\times 3$ log-spaced bins, aligned to the 1D case as close as possible. The 2DPS for the fiducial lightcone (see Fig. 1) is shown in Fig. 3. The panels correspond to the mean redshifts of the lightcone chunks, coinciding with the ones of the 1DPS (see Fig. 2). The columns correspond to the cases with and without noise, as labeled at the top. We can verify that both the cosmic signal and the noise are anisotropic. For the cosmic signal, this is most evident for $z\lesssim 10$ where the noise is subdominant. For the noise, the same can be clearly seen for higher redshifts, where the noise dominates the signal at high $k_{\|}$ much more than in high $k_{\perp}$.
@@ 3.3 Wavelet Scattering Transform
The wavelet transform (Gabor, 1946; Goupillaud et al., 1984; Trott et al., 2012) and wavelet scattering transform (WST) (Mallat, 2011) are based on convolutional filters designed on a Fourier basis with an additional Gaussian envelope along the frequency direction. As such, they capture local features in both image and Fourier domains, which is often relevant in audio and image processing. Higher-order convolutions combined with neural networks (wavelet-based CNNs) have been very successful for such purposes (Bruna & Mallat, 2012; Sifre & Mallat, 2013; Anden & Mallat, 2014). The WST has been extensively explored in cosmological (e.g. Cheng et al. 2020) and 21-cm (e.g. Greig et al. 2022, 2023; Zhao et al. 2023a; Hothi et al. 2023) analyses.
Here we briefly introduce how the WST is defined. It represents a convolution of an input image $I(\bm{x})$ by a set of wavelet (here specifically Morlet) filters. They are characterized by the physical scale $j$ and rotation moment $l$. The actual physical scale of the filter $\psi^{j,l}$ will then be $2^{j}$ pixels, and its orientation angle $l\cdot\pi/L$. The wavelet coefficients correspond to the average of the convolved images, to which we refer as simply wavelets. Note that in this work we restrict ourselves to the 2D WST (see e.g. Zhao et al. 2023a for recent work using the 3D transform).
The first and second order wavelets can be written as:
@@
One can see the similarity between first order wavelets and the PS by writing $\mathrm{PS}(\bm{k})=\langle|I(\bm{x})*\psi^{\prime}|^{2}\rangle$, with $\psi^{\prime}=e^{-i\bm{k}\cdot\bm{x}}$. To extract only cylindrically-averaged information imprinted in the 21-cm signal (analogous to calculating 2DPS in a Fourier basis), one can average coefficients over the rotations:
@@
For further information about the WST and wavelet coefficients we refer the reader to Cheng et al. (2020); Greig et al. (2022, 2023).
Our setup for calculating wavelets is the following. We first select a series of 8 images, centered at the same redshifts as for the power spectra. For each image we calculate the first and second-order wavelets and combine them for the final summary. As the WST scale $2^{J}\times 2^{J}$ has to be smaller than the sky-plane images ($50\times 50$), we set $J=4$. Further, we fix $L=4$ as a common choice (Cheng et al., 2020; Greig et al., 2022). For computing convolutions, we use the scattering_transform [^6] https://github.com/SihaoCheng/scattering_transform package (Cheng et al., 2020). For each redshift slice, we average the first and second order coefficients computed on a rolling 32x32 window. Our final WST summary then consists of $J=4$ $S_{1}$ and $J(J+1)/2=10$ $S_{2}$ coefficients for each of the 8 redshift slices, for a total of 112 numbers.

### v2

@@ 2.2 Finite differencing
To numerically evaluate Eq. 4, we grouped the terms into ?fixed? ($\bm{\mu}$, $\bm{\Sigma}$, $\bm{\Sigma}^{-1}$) and ?differentiated? ($\nabla_{\bm{\theta}}\,\bm{\mu}$, $\nabla_{\bm{\theta}}\,\bm{\Sigma}$). We calculated fixed terms from realizations of the summary around the same parameter value $\bm{\theta}^{*}$. To compute the mean, we can use the maximum likelihood estimator (MLE):
@@
where the index $i$ labels different realizations of initial conditions and other sources of stochasticity (e.g., telescope noise). It is known that such an estimator of the mean is unbiased. However, for the covariance matrix the situation is more complex, as the MLE
@@
is biased. One can show that the estimators of the covariance and its inverse can be unbiased in the following way (e.g., Hartlap et al. 2007):
@@
A similar approach is used to estimate the ?differentiated parameters.? In the case in which the simulator is not differentiable, we can use finite differencing to estimate gradients of the mean and covariance:
@@
Here, $\delta\bm{\theta}_{m}$ represents a small change in the parameter over which the derivative is computed and $\delta\bm{d}(\bm{\theta}^{*})_{i}/\delta\bm{\theta}_{m}$ corresponds to the numerical derivative of a given summary realization (i.e., fixing all sources of stochasticity).
@@ 2.3 The distribution of the Fisher information over parameter space
Below we show examples of common 21 cm summaries evaluated at a fiducial set of parameters, $\bm{\theta}_{\text{fid}}$. Unfortunately, the cosmic 21 cm signal does not yet have a well-defined fiducial model and our choice of $\bm{\theta}_{\text{fid}}$ is only a ?best guess.? Therefore, to properly assess the quality of a particular summary, we sampled many points $\bm{\theta}^{*}$ from the prior $P(\bm{\theta})$, calculating the Fisher information $\det\bm{F}(\bm{\theta}^{*})$ at each sample. We used this prior-weighted distribution of Fisher information as the main metric for assessing the quality of a given summary.666 Note that the prior-weighted distribution of the Fisher information is closely related to the Mutual Information $I(\bm{x},\bm{\theta})=D_{\text{K}L}(p(\bm{x},\bm{\theta})\|p(\bm{x})p(\bm{\theta}))$, where $D_{\text{K}L}$ is Kullback-Liebler divergence (Brunel & Nadal, 1998). Mutual Information can also be used to asses the quality of a summary (e.g., Sui et al. 2023).
3 The cosmic 21 cm signal and its summary statistics
Figure 1: Schematic illustrating our pipeline for producing data summaries for a given astrophysical parameter combination (adapted from Prelogović & Mesinger 2023). Starting from a realization of the cosmological signal simulated with 21cmFAST, we remove the mean at each redshift and add a telescope noise realization corresponding to a 1000h observation with SKA1-Low. The resulting light cone is compressed into different summaries. Explicitly defined summaries are marked in blue, while neural network summaries are marked in red. This realization is computed at the fiducial parameter set, which was used only for IMNN training and visualizations of the summaries. See the main text for more details.
For a given sample of astrophysical parameters, $\bm{\theta}^{*}$, we computed a realization of mock data using the following steps (illustrated in Fig. 1):
@@ 3 The cosmic 21 cm signal and its summary statistics
    Sampling cosmological initial conditions, we simulated a realization of the 21 cm lightcone using the 21cmFAST code.
@@
    We subtracted the mean from each frequency slice, mimicking an interferometric observation.
@@
    We simulated a $1000\mathrm{h}$ tracked scan with the SKA1-low telescope, including the corresponding $uv$ coverage and realization of the instrument noise.
@@
    We compressed the resulting light cone into a given summary observation.
We elaborate on these steps below.
@@
The cosmic 21 cm signal was simulated using the semi-numerical code, 21cmFAST v3777https://github.com/21cmfast/21cmFAST (Mesinger & Furlanetto, 2007; Mesinger et al., 2011; Murray et al., 2020). In brief, we used the Park et al. (2019) galaxy model, varying five parameters that characterize the unknown UV and X-ray properties of high-$z$ galaxies through scaling relations:
@@
  • $\bm{f_{*,10}}$ – the fraction of galactic gas in stars, normalized at the halo mass of $10^{10}M_{\odot}$. The stellar to halo mass relation of the faint galaxies that drive reionization is well described by a power law: $M_{*}/M_{h}=f_{*,10}(M_{h}/10^{10}M_{\odot})^{0.5}(\Omega_{b}/\Omega_{m})$ (see Park et al. 2019 and references therein).
@@
  • $\bm{f_{\text{esc},10}}$ – the ionizing UV escape fraction, normalized at the halo mass $10^{10}M_{\odot}$. Again, the (mean) escape fraction is assumed to be a power law with halo mass, and here we fixed its index to $\alpha_{\text{esc}}=-0.5$ (e.g., Qin et al. in prep).
@@
  • $\bm{M_{\text{turn}}}$ – the characteristic host halo mass below which galaxies are inefficient at forming stars, due to slow gas accretion, supernovae feedback, and/or photo-heating.
@@
  • $\bm{E_{0}}$ – the characteristic X-ray energy below which photons are absorbed by the interstellar medium (ISM) of the host galaxy.
@@
  • $\bm{L}_{\bm{X}<2\mathrm{keV}}/\mathrm{SFR}$ – the X-ray soft band (in the energy range $E_{0}$ – 2 keV) luminosity per star formation rate (SFR), in units of $\mathrm{erg\,s^{-1}\,keV^{-1}\,M_{\odot}^{-1}\,yr}$. The SFR is taken to be: $\dot{M}_{*}=M_{*}/(0.5H(z)^{-1})$, where the Hubble time, $H(z)^{-1}$, also scales with the dynamical time at the virial radius.
@@
Our fiducial, $\bm{\theta}_{\text{fid}}$ parameter vector corresponds to the following: $\log_{10}f_{*,10}=-1.3$, $\log_{10}f_{\text{esc}}=-1$, $\log_{10}M_{\text{turn}}[M_{\odot}]=8.7$, $\log_{10}L_{X<2\mathrm{keV}}/\text{SFR}[\mathrm{erg\,s^{-1}\,keV^{-1}\,M_{\odot}^{-1}\,yr}]=40$, $E_{0}[\mathrm{keV}]=0.5$. This particular choice is only used when training the IMNN, and for the visualizations of the summaries below. More details on the model and its motivations are provided in Park et al. 2019 and references therein.
@@
To simulate the observation, we followed the procedure described in Prelogović et al. (2022); Prelogović & Mesinger (2023) (c.f. Fig. 1). In brief, we calculated the 21 cm signal from redshifts 30 to 5, on a 300 Mpc coeval cube, with 1.5 Mpc resolution. The snapshots of the evolved coeval cube were then stacked to produce the final light cone of the signal. After the subtraction of the mean for every sky-plane slice, we added thermal noise corresponding to a a $6\,\mathrm{h/day}$, $10\,\mathrm{s}$ integration time tracked scan with SKA1-Low antenna configuration for a $1000\,\mathrm{h}$ observation in total. For this task we used tools21cm [^8] https://github.com/sambit-giri/tools21cm, where a separate UV coverage and noise is calculated for each sky-plane slice. Before computing each summary, we additionally smoothed the light cone with a box-car filter obtaining a final resolution of 6 Mpc. This final step does not impact our results as it is comparable to the SKA1-Low beam, but is needed in order to minimize GPU memory usage for certain summaries (see below for more details). We also note that in this work we made the optimistic assumption of perfect foreground removal, thus exploring the maximum future potential of our set of summary statistics.
@@
The resulting 3D light cone was compressed into a given summary statistic. We explore the following summaries in this work:
@@
    – spherically averaged 1D power spectrum,
@@
    – cylindrically averaged 2D power spectrum,
@@
    – 2D wavelet scattering transform,
@@
    – recurrent neural network,
@@
    – information-maximizing neural network.
@@
    – concatenation of the 2DPS and IMNN summaries.
The first four are explicitly defined while the following two are ?learned? by neural networks (NNs). The last, 2DPS + IMNN, is a combination of the two individual summaries. We describe them in detail below.
3.1 Spherically averaged (1D) power spectrum
Figure 2: 1DPS for the fiducial light cone shown in Figure 1. Different colors correspond to different wave modes, while solid (dashed) lines correspond to light cones with (without) noise. Comparing the solid and dashed lines, we see that the signal is noise dominated at higher $k$ modes and redshifts (see also Fig. 1). All power spectra are computed from comoving cubes extracted from the light cone, centered on the redshifts indicated by the vertical dashed lines.
As discussed in the introduction, the 1DPS, is the most common choice of a summary statistic. Despite the cosmic 21 cm signal being non-Gaussian, the 1DPS is a natural choice when seeking to maximize the S/N of an interferometric first detection.
The 1DPS of the 21 cm signal $\delta T_{b}(\bm{x},z)$, where $\bm{x}$ is the sky-plane coordinate and $z$ the redshift, is defined as
@@ 3.1 Spherically averaged (1D) power spectrum
The $\delta T_{b}(\bm{k},z^{\prime})$ is calculated from the Fourier transforms of mean-subtracted light cone segments centered around $z^{\prime}$. Here we calculated the 1DPS in three log-spaced $k$-bins, computed at eight different $z^{\prime}$. In Fig. 2 we show the 1DPS of the fiducial light cone (see Fig. 1). Solid curves correspond to the full signal: cosmic + noise, while the dashed curves are only the cosmic signal. We see the usual three peaked redshift evolution of the large-scale power at $k\sim$1 Mpc-1, tracing fluctuations in the IGM ionized fraction, the IGM temperature and Lyman alpha background (e.g., Pritchard & Furlanetto 2007; Mesinger et al. 2014). By comparing the solid and dashed curves, we see that the signal is noise dominated at small scales and early times. Indeed, $k\geq$ 0.4 Mpc-1 remains noise dominated virtually at all redshifts ($z>7$), for this fiducial model and choice of 1000h integration time.
@@ 3.2 Cylindrical (2D) power spectrum
Figure 3: 2DPS of the fiducial model shown in Figure 1. Different columns show the 2DPS for the pure cosmic signal and including thermal noise, as labeled on the top. Each pair corresponds to a different central redshift, coinciding with those used in the 1DPS summary (see Fig. 2). Furthermore, the bins between 1DPS and 2DPS coincide as much as possible. As in the 1DPS case, high-$k$ modes and high redshifts are noise dominated.
The cylindrically averaged 2DPS distinguishes between sky-plane ($k_{\perp}$) and line-of-sight ($k_{\|}$) modes. Redshift space distortions (e.g., Bharadwaj & Ali 2005; Barkana & Loeb 2005) and light cone evolution (e.g., Greig & Mesinger 2018; Mondal et al. 2018) result in an anisotropic cosmic signal. Therefore, the 2DPS should encode more physical information compared to the 1DPS. Moreover, the instrument/foregrounds are better characterized in the 2DPS (e.g., see the review in Liu et al. 2014b), making it a natural summary observable for preliminary, low S/N measurements.
The 2DPS is defined as
@@
where the Fourier transform is performed on the same light cone chunks as in the 1D case, with the difference being that the expectation value is calculated over $(k_{\perp},k_{\|})$ bins. Here we picked $3\times 3$ log-spaced bins, aligned to the 1D case as close as possible. The 2DPS for the fiducial light cone (see Fig. 1) is shown in Fig. 3. The panels correspond to the mean redshifts of the light cone chunks, coinciding with the ones of the 1DPS (see Fig. 2). The columns correspond to the cases with and without noise, as labeled at the top. We can verify that both the cosmic signal and the noise are anisotropic. For the cosmic signal, this is most evident for $z\lesssim 10$ where the noise is subdominant. For the noise, the same can be clearly seen for higher redshifts, where the noise dominates the signal at high $k_{\|}$ much more than in high $k_{\perp}$.
@@ 3.3 Wavelet scattering transform
The wavelet transform (Gabor, 1946; Goupillaud et al., 1984; Trott et al., 2012) and WST (Mallat, 2011) are based on convolutional filters designed on a Fourier basis with an additional Gaussian envelope along the frequency direction. As such, they capture local features in both image and Fourier domains, which is often relevant in audio and image processing. Higher-order convolutions combined with NNs (wavelet-based convolutional NNs) have been very successful for such purposes (Bruna & Mallat, 2012; Sifre & Mallat, 2013; Anden & Mallat, 2014). The WST has been extensively explored in cosmological (e.g., Cheng et al. 2020) and 21 cm (e.g., Greig et al. 2022, 2023; Zhao et al. 2023a; Hothi et al. 2023) analyses.
Here we briefly introduce how the WST is defined. It represents a convolution of an input image $I(\bm{x})$ by a set of wavelet (here specifically Morlet) filters. They are characterized by the physical scale $j$ and rotation moment $l$. The actual physical scale of the filter $\psi^{j,l}$ will then be $2^{j}$ pixels, and its orientation angle $l\cdot\pi/L$. The wavelet coefficients correspond to the average of the convolved images, to which we refer as simply wavelets. We note that in this work we restricted ourselves to the 2D WST (see, e.g., Zhao et al. 2023a for recent work using the 3D transform).
The first- and second-order wavelets can be written as
@@
One can see the similarity between first-order wavelets and the PS by writing $\mathrm{PS}(\bm{k})=\langle|I(\bm{x})*\psi^{\prime}|^{2}\rangle$, with $\psi^{\prime}=e^{-i\bm{k}\cdot\bm{x}}$. To extract only cylindrically averaged information imprinted in the 21 cm signal (analogous to calculating 2DPS in a Fourier basis), one can average coefficients over the rotations:
@@
It is important to mention that this procedure removes anisotropic information. Alternative definitions such as the reduced WST (Allys et al., 2019) are able to reduce the coefficient dimensionality, while keeping anisotropy. For further information about the WST and wavelet coefficients we refer the reader to Cheng et al. (2020) and Greig et al. (2022, 2023).
Our setup for calculating wavelets is the following. We first selected a series of eight images, centered at the same redshifts as for the power spectra. For each image we calculated the first- and second-order wavelets and combined them for the final summary. As the WST scale ($2^{J}\times 2^{J}$) has to be smaller than the sky-plane images ($50\times 50$), we set $J=4$. Furthermore, we fixed $L=4$ (Cheng et al., 2020; Greig et al., 2022). To compute convolutions, we used the scattering_transform [^9] https://github.com/SihaoCheng/scattering_transform package (Cheng et al., 2020). For each redshift slice, we averaged the first- and second-order coefficients computed on a rolling 32x32 window. Our final WST summary thus consists of $J=4$ $S_{1}$ and $J(J+1)/2=10$ $S_{2}$ coefficients for each of the eight redshift slices, for a total of 112 numbers.
Figure 4: First-order wavelet coefficients (Eqs. 14 and 16) for the fiducial light cone shown in Figure 1. Different colors correspond to different $J$ modes, while solid (dashed) lines correspond to the case with (without) noise. Comparing the solid and dashed lines, we see that the signal is noise dominated at smaller scales (smaller $J$) and higher redshifts (see also Fig. 1).
In Figure 4, we show the redshift evolution of first-order wavelet coefficients ($S_{1}$; see Eqs. 14, 16). Solid (dashed) lines correspond to coefficients with (without) telescope noise. We see the same qualitative trends as in Fig. 2, with lower $J$ modes (i.e., smaller scales) more noise dominated. As a result, these low $J$ modes contribute less to the total Fisher information (Hothi et al., 2023).